OpenStudy (anonymous):

I need to evaluate (sin5theta)/(theta +(tan8theta))

OpenStudy (anonymous):

\[\frac{\sin(5\theta)}{\theta+\tan(8\theta)}\] largely depends on \(\theta\)

OpenStudy (anonymous):

sorry, I forgot to type in the whole question. I thought if I pressed enter it would double space, but instead it just submitted it! I need to evaluate the limit as theta apporaches 0. I get that it simplifies to \[[\lim \theta \rightarrow 0 (\sin 5\theta/\theta)]/[\lim \rightarrow0(\sin8\theta/\theta) \times \lim \theta \rightarrow0(1/\cos8\theta)\], but this is where I get lost on how to evaluate it. I know the what to do when it's sin(theta)/(theta) but not when there's an eight or five infront of it. I thought it would just simplify down to zero as well since (5x0)=0, but it does not. 1/2 is wrong and so is 5/2. Please help!!!

OpenStudy (anonymous):

I again typed the equation wrong, I know there's a one plus the limit of sin8theta/theta. I included that in my attempting to answer, I just forgot to type it.